OpenStudy (brixzii):

If m varies directly as the square root of y, inversely as p2, and directly as n, what happens to m when y is quadrupled, p is tripled, and n is multiplied by 5?

OpenStudy (danjs):

\[m ~~\alpha ~~\frac{ \sqrt{y} * n}{ p^2 }\]

OpenStudy (danjs):

put in those numbers on each variable.. \[\frac{ \sqrt{4y} * 5n}{ (3p)^2 } \]

OpenStudy (danjs):

work out what those numbers simplify to, so you will have some # times the original expression...